OpenStudy (anonymous):

Simplify: (4x)2(-2y3) -8x2y3 None of the above -32x2y3 -8xy6 2. Simplify: (3x+2y)(3x-2y)= 9x2+12xy+4y2 None of the above 9x2+4y2 9x2-4y2

OpenStudy (mathstudent55):

\((4x)^2(-2y^3) \) What does squaring a quantity mean?

OpenStudy (anonymous):

1=none of the above what is 2

OpenStudy (mathstudent55):

1 is incorrect

OpenStudy (anonymous):

then what would it be

OpenStudy (mathstudent55):

Would you like to answer my question above? We'll go over the problem step by step.

OpenStudy (mathstudent55):

The left side of the first problem is squaring the quantity 4x. What does squaring a quantity mean?

OpenStudy (mathstudent55):

Here's a simpler question. What do you do to calculate \(5^2\) ?

OpenStudy (anonymous):

25

OpenStudy (mathstudent55):

You gave me the answer, and the answer is correct, but I asked what do you do to find the answer? You multiply 5 by 5, right?

OpenStudy (anonymous):

ya man

OpenStudy (mathstudent55):

In your first problem, you need to square 4x. What do you do to square 4x? You must multiply 4x by 4x. What is 4x * 4x?

OpenStudy (anonymous):

16

OpenStudy (mathstudent55):

The 4 times the 4 equals 16. That's good, but you also have x times x. \((4x)^2 = 4x \times 4x = 16x^2 \)

OpenStudy (mathstudent55):

The entire problem is: \((4x)^2(-2y3) \) \(= 16x^2(-2y^3) \) \(=-32x^2y^3\) Notice that one of the choices above is the answer.

OpenStudy (anonymous):

how bout this problem Simplify: 9x^3+6x^2+3x/3x None of the above 3x2+2x+1 3x2-2xy-2 3x4+6x3+x2

OpenStudy (mathstudent55):

Wait, didn't you want to do the second problem above also?

OpenStudy (mathstudent55):

2. Simplify: (3x+2y)(3x-2y)= 9x2+12xy+4y2 None of the above 9x2+4y2 9x2-4y2

OpenStudy (mathstudent55):

This is a famous multiplication in math. When you multiply a sum and a difference of the sdame two quantities. The general pattern is: \( (a + b)(a - b) = a^2 - b^2\) Notice that the two things you are adding, a and b, you are also subtracting, a - b. You are multiplying a sum and a subtraction (difference) off the same two things. All you need to do is square both of the things and subtract them. a^2 - b^2.

OpenStudy (mathstudent55):

Look at your problem: (3x+2y)(3x-2y)= You are adding 3x and 2y, and you are subtracting 3x - 2y. To multiply them together, square the two things, (3x)^2 and (2y)^2 and subtract them. \( (3x + 2y)(3x - 2y) = \) \(= (3x)^2 - (2y)^2 \) \(= 9x^2 - 4y^2 \) This is one of your choices up there.

OpenStudy (mathstudent55):

Simplify: \( \dfrac{9x^3+6x^2+3x}{3x} \) Divide every term in the numerator by the denominator. \( \dfrac{9x^3+6x^2+3x}{3x} \) \( = \dfrac{9x^3}{3x} + \dfrac{6x^2}{3x} + \dfrac{3x}{3x} \) \(= 3x^2 + 2x + 1\)

OpenStudy (anonymous):

Simplify: (7x-2)+(2x+17)= 9x+19 9x2+19 9x+15 9x2+15